# An Efficient and Precise Training Data Construction Framework for Process-supervised Reward Model in Mathematical Reasoning

### v1

@@ Abstract
Enhancing the mathematical reasoning capabilities of Large Language Models (LLMs) is of great scientific and practical significance. Researchers typically employ process-supervised reward models (PRMs) to guide the reasoning process, effectively improving the models’ reasoning abilities. However, existing methods for constructing process supervision training data, such as manual annotation and per-step Monte Carlo estimation, are often costly or suffer from poor quality. To address these challenges, this paper introduces a framework called EpicPRM (Efficient, Precise, Cheap), which annotates each intermediate reasoning step based on its quantified contribution and uses an adaptive binary search algorithm to enhance both annotation precision and efficiency. Using this approach, we efficiently construct a high-quality process supervision training dataset named Epic50k, consisting of 50k annotated intermediate steps. Compared to other publicly available datasets, the PRM trained on Epic50k demonstrates significantly superior performance. Getting Epic50k at  [^1] https://github.com/xiaolizh1/EpicPRM.
@@
Wei Sun1,2, Qianlong Du1, Fuwei Cui1  and Jiajun Zhang1,2,3 1Institute of Automation, Chinese Academy of Sciences 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3Wuhan AI Search {sunwei2023,fuwei.cui}@ia.ac.cn {qianlong.du,jjzhang}@nlpr.ia.ac.cn
Figure 1: An examples of getting the correct answer based on the wrong steps. On the basis of $s_{6}$, completer continue to sample a number of rollouts. We can see that since the completer has a certain error correction capability, even if there is an obvious error in step 6.
@@ 1 Introduction
Improving the mathematical reasoning ability of Large Language Models (LLMs) can significantly accelerate advancements in artificial intelligence across various scientific domains. To enhance the mathematical reasoning capability of LLMs, several studies Yu et al. (2023a); Cobbe et al. (2021); Shao et al. (2024) employ Outcome-supervised Reward Models (ORMs) to guide Chain-of-Thought (CoT) generation and leverage reinforcement learning to optimize generative performance. Among these approaches, process supervision, leveraging Process-supervised Reward Models (PRMs), has emerged as a more effective approach Li et al. (2022); Uesato et al. (2022); Lightman et al. (2023); Chen et al. (2024b), providing substantial improvements through step-by-step supervision.
@@ 3.1.1 Perplexity Instead of Count
Using MC estimation to represent the probability of obtaining the correct answer from a given state has inherent limitations. When $N$ is not large enough, counting will be accidental, we argue that using $M/N$ as a direct representation of the probability of sampling correct rollouts is imprecise. Therefore, we propose using perplexity (PPL) calculated by logarithmic probability to represent the sampling probability of each rollout:
@@ 4.6 Out of Domain Generalization
To evaluate our PRM on stronger generators and out of domain benchmarks, we conducted experiments on PROCESSBENCH Zheng et al. (2024), a PRM assessment benchmark covering GSM8K, MATH, OlympiadBench, and Omni-MATH. Solutions were generated by LLaMA and Qwen models, including Qwen2.5-Math-7B/72B-Instruct, and labeled step-wise by human experts. Results are shown in Table 3. Epic50k demonstrates strong generalization across benchmarks of varying difficulty. While it lags behind human-labeled PRM800k, it performs competitively with open-source PRMs, excelling at Olympiad-level problems where it matches Skywork-PRM-7B. Though Qwen2-1.5B-Epic50k scores slightly below Skywork-PRM-1.5B (by <1%). Skywork leverages Qwen2.5-MATH-1.5B-Instruct, whereas our Qwen2-1.5B-Epic50k is based on the previous-generation Qwen2-MATH-1.5B-Base. Given the disparity in base models, this advantage is expected. Notably, despite being trained only on MATH, Qwen2-1.5B-Epic50k surpasses Skywork-PRM-1.5B on OlympiadBench and Omni-MATH, highlighting strong generalization from easier to harder problems.
@@ 5 Related Work
Prompt-based approaches. Particularly Chain-of-Thought (CoT) Wei et al. (2022), have significantly advanced mathematical reasoning Perez et al. (2021). Subsequent work has explored CoT fine-tuning Ouyang et al. (2022); Li et al. (2024); Mitra et al. (2024); Yu et al. (2023b) and extended prompt engineering to problem decomposition Zhou et al. (2022); Hao et al. (2023) and programming Chen et al. (2022); Li et al. (2023); Zhou et al. (2023); Chen et al. (2024a); Yang et al. (2024b). Selecting optimal solutions from multiple sampled paths has proven more effective than prompt engineering alone. Self-Consistency Wang et al. (2022) uses majority voting across generated CoT trajectories, while RAP Hao et al. (2023) uses self-evaluation and feedback for sequential subproblem solving. However, existing research Huang et al. (2023) indicates that models encounter significant difficulty in self-assessing without external information.

### v2

@@ Abstract
Enhancing the mathematical reasoning capabilities of Large Language Models (LLMs) is of great scientific and practical significance. Researchers typically employ process-supervised reward models (PRMs) to guide the reasoning process, effectively improving the models’ reasoning abilities. However, existing methods for constructing process supervision training data, such as manual annotation and per-step Monte Carlo estimation, are often costly or suffer from poor quality. To address these challenges, this paper introduces a framework called EpicPRM (Efficient, Precise, Cheap), which annotates each intermediate reasoning step based on its quantified contribution and uses an adaptive binary search algorithm to enhance both annotation precision and efficiency. Using this approach, we efficiently construct a high-quality process supervision training dataset named Epic50k, consisting of 50k annotated intermediate steps. Compared to other publicly available datasets, the PRM trained on Epic50k demonstrates significantly superior performance. Getting Epic50k at https://github.com/xiaolizh1/EpicPRM.
@@
Wei Sun1,2, Qianlong Du1, Fuwei Cui1  and Jiajun Zhang1,2,3 1Institute of Automation, Chinese Academy of Sciences 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3Wuhan AI Research {sunwei2023,fuwei.cui}@ia.ac.cn {qianlong.du,jjzhang}@nlpr.ia.ac.cn
footnotetext: Corresponding author.
Figure 1: An examples of getting the correct answer based on the wrong steps. On the basis of $s_{6}$, completer continues to sample a number of rollouts. We can see that since the completer has a certain error correction capability, even if there is an obvious error in step 6.
@@ 1 Introduction
Improving the mathematical reasoning ability of Large Language Models (LLMs) can significantly accelerate advancements in artificial intelligence across various scientific domains. To enhance the mathematical reasoning capability of LLMs, several studies Yu et al. (2023a); Cobbe et al. (2021); Shao et al. (2024); Chen et al. (2025a) employ Outcome-supervised Reward Models (ORMs) to guide Chain-of-Thought (CoT) generation and leverage reinforcement learning to optimize generative performance. Among these approaches, process supervision, leveraging Process-supervised Reward Models (PRMs), has emerged as a more effective approach Li et al. (2022); Uesato et al. (2022); Lightman et al. (2023); Chen et al. (2024b), providing substantial improvements through step-by-step supervision.
@@ 3.1.1 Perplexity Instead of Count
Using MC estimation to represent the probability of obtaining the correct answer from a given state has inherent limitations. When $N$ is not large enough, counting will be accidental, we argue that using $M/N$ as a direct representation of the probability of sampling correct rollouts is imprecise. According to the Law of Large Numbers, estimating probabilities using counting methods necessitates a large number of samples to achieve reliable accuracy. For instance, observing two heads in two coin tosses does not imply a 100% probability of obtaining heads. Similarly, if we sample $N$ rollouts and find $M$ correct ones, estimating the probability of correctness as $M/N$ inherently carries uncertainty due to sampling variability. By using perplexity, we can directly calculate the probability of the model generating each rollout, thereby mitigating errors introduced by sampling. Furthermore, if a correct rollout has an exceptionally low generation probability but is included in the sample, a counting-based method would overestimate the likelihood of obtaining correct rollouts. Therefore, we propose using perplexity (PPL) calculated by logarithmic probability to represent the sampling probability of each rollout:
@@ 4.5.1 Compared Search Algorithms
Qwen2.5-1.5B-Instruct-PRM800k 49.1 59.3 32.5 35.9 44.2
Qwen2.5-1.5B-Instruct-Shepherd 44.1 29.8 12.1 17.6 25.9
Qwen2.5-1.5B-Instruct-Epic50k 58.1 53.2 32.2 40.6 46.0
@@ 4.6 Out of Domain Generalization
To evaluate our PRM on stronger generators and out of domain benchmarks, we conducted experiments on PROCESSBENCH Zheng et al. (2024), a PRM assessment benchmark covering GSM8K, MATH, OlympiadBench, and Omni-MATH. Solutions were generated by LLaMA and Qwen models, including Qwen2.5-Math-7B/72B-Instruct, and labeled step-wise by human experts. Results are shown in Table 3. Epic50k demonstrates strong generalization across benchmarks of varying difficulty. While it lags behind human-labeled PRM800k, it performs competitively with open-source PRMs, excelling at Olympiad-level problems where it matches Skywork-PRM-7B. Though Qwen2-1.5B-Epic50k scores slightly below Skywork-PRM-1.5B (by <1%). Skywork leverages Qwen2.5-MATH-1.5B-Instruct, whereas our Qwen2-1.5B-Epic50k is based on the previous-generation Qwen2-MATH-1.5B-Base. Given the disparity in base models, this advantage is expected. Notably, despite being trained only on MATH, Qwen2-1.5B-Epic50k surpasses Skywork-PRM-1.5B on OlympiadBench and Omni-MATH, highlighting strong generalization from easier to harder problems. In order to be consistent with the base model of Skywork-PRM-1.5B, we conducted comparative experiments based on the Qwen2.5-1.5B-Instruct model. Our Epic50k performs significantly better than other methods.
@@ 5 Related Work
Improving the mathematical reasoning ability of Large Language Models (LLMs) can significantly accelerate advancements in artificial intelligence across various scientific domains, such as Translation He et al. (2025), Scientific agent reasoning Ren et al. (2025), Long-text reasoning Wan et al. (2025); Chen et al. (2025b, a, c), Multilingual Ghosh et al. (2025); Yang et al. (2023), Multi-modal Huang et al. (2025); Jian et al. (2024). Prompt-based approaches. Particularly Chain-of-Thought (CoT) Wei et al. (2022), have significantly advanced mathematical reasoning Perez et al. (2021). Subsequent work has explored CoT fine-tuning Ouyang et al. (2022); Li et al. (2024); Mitra et al. (2024); Yu et al. (2023b) and extended prompt engineering to problem decomposition Zhou et al. (2022); Hao et al. (2023) and programming Chen et al. (2022); Li et al. (2023); Zhou et al. (2023); Chen et al. (2024a); Yang et al. (2024b). Selecting optimal solutions from multiple sampled paths has proven more effective than prompt engineering alone. Self-Consistency Wang et al. (2022) uses majority voting across generated CoT trajectories, while RAP Hao et al. (2023) uses self-evaluation and feedback for sequential subproblem solving. However, existing research Huang et al. (2023) indicates that models encounter significant difficulty in self-assessing without external information.
@@ 6 Conclusion
Acknowledgements
We thank our colleagues Jianghao Chen, Jian Pu, Wen Yang, Cong Li, Tengxiao Xi, Tianyu Peng, Ziliang Pang, Junhong Wu for their insightful and constructive feedback. We thank Qian Li and Zhenggang Piao for their special assistance. In addition, we thank all reviewers for their valuable comments and recognition of our work. This work is supported by National Key R&\&&D Program of China N0.2022ZD0160602 and the Strategic Priority Research Program of Chinese Academy of Sciences under Grant N0.XDA04080400.